a 50 ml sample of .1 mol/L HCIO(aq) is titrated with .2 mol/L KOH(aq). find ph given the following volumes of added base: (Ka= 2.8 x 10-8) a) 0.0 ml b) 10.0 ml c) 25.0 ml d) 30.0 ml. show your work


1
Expert's answer
2022-01-12T07:06:01-0500



CH"_3" COOH "+" NaOH "\to" CH"_3" COONa"+H_2O"



Number of moles of CH"_3" COOH

"=0.150×\frac{50}{1000}=0.0075\>moles"


Number of moles of NaOH

"=0.150×\frac{20}{1000}"

"0.003\>moles"



Moles of excess acid"=0.0075-0.003"

"=0.0045\>moles"



PH"=" "pk_a+" log "\frac{[A^-]}{[HA]}"

"=-log\>1.8×10^{-5}+log\>\frac{0.003}{0.0045}"


"=4.7447-0.1761"


"=4.586"